the second angle of a triangle is 30 degrees less than the first angle. the third angle is twice the second angle. how large are the angles?
please and thank you

Answer by orca(409)

let x be the measure of the first angle.
Then the second angle is x - 30, and the third angle is 2(x-30).
Their sum is x + (x - 30) + 2(x - 30)
As the three angles of any triangle add up to 180 degrees, we have:
x + (x - 30) + 2(x - 30)=180
Solving for x, we have:
x + x - 30 + 2x - 60 = 180
4x -90 = 180
4x = 270
x = 67.5
So
The first angle is 67.5 degrees
The second angle is x-30 = 67.5-30=37.5 degrees
The third is twice the second. so it is 2*37.5 = 75 degrees